Consider the demand function for processed pork in Canada,

Qd=386.00-20p + 20pb + 3pc + 0.002Y

The supply function for processed pork in Canada is:

Qs=240.00 + 30p - 60ph

pb is the price of beef $4 per kg

pc is the price of chicken $3 per kg

Y is the income of consumers $12,500

ph is the price of a hog $1.50 per kg

p is the price of pork

Q is the quantity of pork demanded (measured in millions of kg per year)

Solve for the equilibrium price and quantity for pork.
Answer #1

Qd = 386 - 20p + 20(4) + 3(3) + 0.002(12500)

Qd = 386 - 20p + 80 + 9 + 25

Qd = 500 - 20p

Qs = 240 + 30p - 60(1.50)

Qs = 150 + 30p

Equilibrium condition;

Qd = Qs

500 - 20p = 150 + 30p

500 - 150 = 30p + 20p

350 = 50p

p = 7

Q = 150 + 30(7)

Q = 150 + 210

Q = 360 units